# Synthetic Modelling of Polarized Dust Emission in Intermediate-Mass YSOs: I: Constraining the Role of Iron Inclusions and Inelastic Relaxation on Grain Alignment with ALMA Polarization

### v2

@@ 3.3 Modelling of grain alignment
(c): Disruption range is examined again with the alignment range.
(d): Fraction of disrupted grains $f_{\rm disr}$ depends on the values of $f_{\rm high-J}$.
@@ 7.7 Impact of stellar radiation field strength on dust polarization
As argued in Section 3.1, we use the high bolometric luminosity of $100L_{\odot}$ to maximize the impact of stellar radiation on grain alignment and dust polarization. This value is beyond the average luminosity of the protostar samples used for comparison in Section 6. However, some of these protostars are intermediate-mass Class 0/Is objects, and these type of objects can have very high bolometric luminosity of $\sim 100L_{\odot}$ (i.e., Serpens SMM1, Le Gouellec et al. 2019, HOPS 288 and 361, Furlan et al. 2016, Cheng et al. 2022). Moreover, the high mass-accretion rate $\dot{M}_{\rm acc}\sim 4\times 10^{-6}M_{\odot}/\rm yr$ corresponding to $L_{\rm center}=100L_{\odot}$ still can happen during the accretion burst  [^9] We note our MHD simulation is for intermediate-mass star formation, with the accretion rate onto the sink particle of $\dot{M}_{\rm acc}\sim 10^{-5}-10^{-4}M_{\odot}/\rm yr$, Le Gouellec et al. (2023b) (i.e., HH34 IRS, Antoniucci et al. 2008, see also modeling of accretion activities by Elbakyan et al. 2016). Therefore, it is acceptable to adopt $L_{\rm center}=100L_{\odot}$ for understanding the grain alignment dynamic and dust polarization properties inside intermediate-mass cores. However, to better understand these activities inside protostars of less-intense irradiation, we consider two other cases of $L_{\rm center}=20L_{\odot}$ ($\dot{M}_{\rm acc}\sim 8.38\times 10^{-7}M_{\odot}/\rm yr$) and $L_{\rm center}=5L_{\odot}$ ($\dot{M}_{\rm acc}\sim 2\times 10^{-7}M_{\odot}/\rm yr$). This simulation was already done in Le Gouellec et al. (2023b) for the ideal RAT alignment model. However, given the strong correlation of the grain rotational velocity and the IA shown in Figures 4 and 4, the effect of $L_{\rm center}$ on the net alignment degree need to be re-examined. We explore this issue in Appendix D, considering two cases of $a_{\rm max}=50\,{\mu\rm{m}}$ and $a_{\rm max}=10\,{\mu\rm{m}}$.
@@
We find that by decreasing $L_{\rm center}$, apart from the reduction of the alignment size range (as found in Le Gouellec et al. 2023b), aligned dust grains still have efficient IA through inelastic relaxation, and further experience stronger MRAT owing to the enhanced magnetic relaxation resulting from the decreased dust temperature (Figures 16 and 16). In terms of the impact of $L_{\rm center}$ on $p-N_{\rm H_{2}}/N_{\rm H_{2},max}$, we found that decreasing $L_{\rm center}$ only induces a clear reduction of the observed polarization degree (as expected in Le Gouellec et al. 2023b) when $a_{\rm max}$ inside the core reduces to below $\leq 10\,{\mu\rm{m}}$ (Figure 17, right panel). Our MRAT alignment model for SPM grains with $N_{\rm cl}=10^{3}$ and $\mu Q=3\times 10^{9}\,{\rm{erg}}\,{\rm{cm}}^{-3}$ fails to explain ALMA observations inside the inner envelope when the central luminosity reduces to below $\leq 5L_{\odot}$(Figure 16, first row). If grains have higher inelasticity, the situation worsens as models with $L_{\rm center}=20L_{\odot}$ cannot reproduce the polarization degree range observed in low/intermediate Class 0 YSOs.
@@ Appendix B Upper limit of iron inclusions inside SPM grains
with $\tau_{\rm 0}=10^{9}$s and $V_{\rm cl}$ the volume of iron cluster including $N_{\rm cl}$ iron atoms, which is given by:
@@
Metallic iron $1.35\times 10^{5}$ 7.87 1 $6.2\times 10^{4}-1.34\times 10^{5}$ $1.6\times 10^{5}-3.6\times 10^{5}$ $8.3\times 10^{5}-2\times 10^{6}$
$\rm Fe_{\rm 3}\rm O_{\rm 4}$ $6.1\times 10^{5}$ 5.17 0.72 $6.5\times 10^{3}-1.3\times 10^{4}$ $1.7\times 10^{4}-3.8\times 10^{4}$ $8.7\times 10^{4}-2.2\times 10^{5}$
$\gamma-\rm Fe_{\rm 2}\rm O_{\rm 3}$ $1.9\times 10^{5}$ 4.9 0.69 $1.9\times 10^{4}-3.8\times 10^{4}$ $5.1\times 10^{4}-1.1\times 10^{5}$ $2.5\times 10^{5}-6.25\times 10^{5}$
@@
Considering the age of our YSO of $\tau_{\rm YSOs}=38.52$ kry (Section 2) [^13] Note that the typical timescale for Class 0 YSOs is $\sim 10^{4}$ yrs and for Class I YSOs if $\sim 10^{5}$ yrs. Given longer considered timescales, bigger iron clusters can have more time to reorient their magnetic moment to the applied magnetic field direction, which slightly increases the boundary for defining SPM., the maximum cluster volume that Brownian-like motions of clusters can overcome the anisotropy barrier energy to produce the net grain magnetic alignment is given by:
@@
One can see that for both metallic iron, $\rm Fe_{\rm 3}\rm O_{\rm 4}$, and $\gamma-\rm Fe_{\rm 2}\rm O_{\rm 3}$, adopting $N_{\rm cl}=10^{4}$ in the rest of the paper still be under the allowance range for SPM grains regardless of their position inside YSOs (except the case of $\rm Fe_{\rm 3}\rm O_{\rm 4}$ inside the envelope with $T_{\rm d}<15$ K-the case of $L_{\rm star}=5L_{\odot}$). Theoretically, we can increase $N_{\rm cl}$ to nearly $10^{5}$ to enhance the grain magnetic alignment of SPM grains inside the disk. However, given the unclear form of iron inclusions and mechanisms producing the big iron cluster inside large grains, it is very risky to adopt $N_{\rm cl}\sim 10^{5}$ to study the grain alignment dynamic inside protostellar/protoplanetary disk environments.
@@ C.1 Whether higher $\phi_{\rm sp}$ helps to enhance the grain alignment inside the disk?
Appendix D Effect of stellar luminosity on polarized dust emission
D.1 On the alignment efficiency
@@
We provide the distribution map of grain alignment size in case $a_{\rm max}=10\,{\mu\rm{m}}$ in Figure 16. Similar to Figure 16, almost aligned dust grains inside protostars with low $L_{\rm center}=5L_{\odot}$ still can achieve efficient IA and magnetic alignment as grains inside objects with high $L_{\rm center}=100L_{\odot}$. But given smaller $a_{\rm max}=10\,{\mu\rm{m}}$, the increase of $a_{\rm align}$ with decreasing $L_{\rm center}$ narrowers the alignment range clearer, inducing lower net observed polarization fraction shown in Figure 17, right panel).
D.2 On the polarization degree distribution
We show in Figure 17 the effect of $L_{\rm center}$ on the variation of polarization degree with normalized gas column density. ALMA observations are overplotted to compare with synthetic results. We consider model RATA--INELASTIC for SPM grains with $N_{\rm cl}=10^{3}$, $\mu Q=3\times 10^{9}\,{\rm{erg}}\,{\rm{cm}}^{-3}$, and $\Theta=90^{\circ}$. The left one shows results with $a_{\rm max}=50\,{\mu\rm{m}}$, and the right one corresponds to the model with $a_{\rm max}=10\,{\mu\rm{m}}$.
One can see that when VLGs up to $50\,{\mu\rm{m}}$ can distribute over the envelope and inner envelope, $L_{\rm center}$ is not an important factor controlling the polarization fraction inside Class 0 protostars. Indeed, the alignment size range around protostars with lower luminosity is narrower owing to the reduced RAT efficiency, from $[\sim 0.8,50\,{\mu\rm{m}}]$ for $L_{\rm center}=100L_{\odot}$ to $[\sim 3-50\,{\mu\rm{m}}]$ for $L_{\rm center}=5L_{\odot}$ (Figure 16, second column). However, since all aligned dust grains still have efficient IA by inelastic relaxation and experience stronger MRAT mechanism (Figure 2), the narrower of the alignment size range is not significant enough to cause the clear reduction of the net alignment efficiency and the observed polarization degree. In contrast, if $a_{\rm max}$ is reduced to $10\,{\mu\rm{m}}$ (right panel), the decrease of alignment size range (from $[\sim 0.8,10\,{\mu\rm{m}}]$ for $L_{\rm center}=100L_{\odot}$ to $[\sim 3-10\,{\mu\rm{m}}]$ for $L_{\rm center}=5L_{\odot}$) reduces the net grain alignment degree inside the core clearer. Consequently, one gets much lower polarization degree around protostars with low $L_{\rm center}$ (similar as finding in Le Gouellec et al. 2023b). Particularly, the RATA--INELASTIC model fails to explain ALMA observations inside the inner envelope and disk if the central luminosity reduces to $<5L_{\odot}$. If dust grains have higher inelasticity (thus experiencing weaker inelastic relaxation), the reduced IA of aligned dust grains inactivates the MRAT alignment model even if $L_{\rm center}=20L_{\odot}$. For a short summary, the bolometric luminosity only deactivates RAT/MRAT mechanism in protostellar cores if the maximum grain size inside the core is below $10\,{\mu\rm{m}}$.

### v3

@@ 3.1 Radiation fields and dust model
Polarization fraction $p(\%)$
Magnetic moment of iron atom $p$
@@ 7.7 Impact of stellar radiation field strength on dust polarization
As argued in Section 3.1, we use the high bolometric luminosity of $100L_{\odot}$ to maximize the impact of stellar radiation on grain alignment and dust polarization. This value is beyond the average luminosity of the protostar samples used for comparison in Section 6. However, some of these protostars are intermediate-mass Class 0/Is objects, and these type of objects can have very high bolometric luminosity of $\sim 100L_{\odot}$ (i.e., Serpens SMM1, Le Gouellec et al. 2019, HOPS 288 and 361, Furlan et al. 2016, Cheng et al. 2022). Moreover, the high mass-accretion rate $\dot{M}_{\rm acc}\sim 4\times 10^{-6}M_{\odot}/\rm yr$ corresponding to $L_{\rm center}=100L_{\odot}$ still can happen during the accretion burst  [^9] We note our MHD simulation is for intermediate-mass star formation, with the accretion rate onto the sink particle of $\dot{M}_{\rm acc}\sim 10^{-5}-10^{-4}M_{\odot}/\rm yr$, Le Gouellec et al. (2023b) (i.e., HH34 IRS, Antoniucci et al. 2008, see also modeling of accretion activities by Elbakyan et al. 2016). Therefore, it is acceptable to adopt $L_{\rm center}=100L_{\odot}$ for understanding the grain alignment dynamic and dust polarization properties inside intermediate-mass cores. However, to better understand these activities inside protostars of less-intense irradiation, we consider two other cases of $L_{\rm center}=20L_{\odot}$ ($\dot{M}_{\rm acc}\sim 8.38\times 10^{-7}M_{\odot}/\rm yr$) and $L_{\rm center}=5L_{\odot}$ ($\dot{M}_{\rm acc}\sim 2\times 10^{-7}M_{\odot}/\rm yr$). This simulation was already done in Le Gouellec et al. (2023b) for the ideal RAT alignment model. However, given the strong correlation of the grain rotational velocity and the IA shown in Figures 4 and 4, the effect of $L_{\rm center}$ on the net alignment degree need to be re-examined. The grain alignment state inside fainter protostars is shown in Appendix D, considering the case of SPM grains with $N_{\rm cl}=10^{3}$ and $\mu Q=3\times 10^{9}\,{\rm{erg}}\,{\rm{cm}}^{-3}$. We find that by decreasing $L_{\rm center}$, apart from the reduction of the alignment size range as increasing the minimum alignment size range (as found in Le Gouellec et al. 2023b), aligned dust grains still have efficient IA through inelastic relaxation, and further experience stronger MRAT owing to the enhanced magnetic relaxation resulting from the decreased dust temperature.
@@
We show in Figure 11 the effect of $L_{\rm center}$ on the relation $p-N_{\rm H_{2}}/N_{\rm H_{2},max}$, considering $L_{\rm center}=5L_{\odot},20L_{\odot},100L_{\odot}$. We consider model RATA--INELASTIC for SPM grains with $N_{\rm cl}=10^{3},\mu Q=3\times 10^{9}\,{\rm{erg}}\,{\rm{cm}}^{-3}$, the observed angle of $\Theta=90^{\circ}$, with the left showing results with $a_{\rm max}=50\,{\mu\rm{m}}$ and the right corresponding to results with $a_{\rm max}.=10\,{\mu\rm{m}}$. One can see that if VLGs up to $50\,{\mu\rm{m}}$ can be distributed over thousands au core scale (left panel), the central luminosity does not affect the observed polarization fraction due to the insignificant change in the alignment size range when varying $L_{\rm center}$ (Figures 18). In contrast, if $a_{\rm max}$ is reduced to below $10\,{\mu\rm{m}}$ (right panel), the increase in the minimum alignment size around low-luminous protostar significantly narrowers the alignment size range (Figure 18), which clearly induces the lower observed polarization fraction (as expected in Le Gouellec et al. 2023b). Our MRAT alignment model for SPM grains with $N_{\rm cl}=10^{3}$ and $\mu Q=3\times 10^{9}\,{\rm{erg}}\,{\rm{cm}}^{-3}$ fails to explain ALMA observations inside the inner envelope when the central luminosity reduces to below $\leq 5L_{\odot}$(Figure 18, first row). If grains have higher inelasticity, the situation worsens as models with $L_{\rm center}=20L_{\odot}$ cannot reproduce the polarization degree range observed in low/intermediate Class 0 YSOs.
@@ Appendix B Upper limit of iron inclusions inside SPM grains
with $\tau_{\rm 0}=10^{-9}$s and $V_{\rm cl}$ the volume of iron cluster including $N_{\rm cl}$ iron atoms, which is given by:
@@
Metallic iron $1.35\times 10^{5}$ 7.87 1 $6.25\times 10^{4}-1.25\times 10^{5}$ $1.67\times 10^{5}-3.75\times 10^{5}$ $8.34\times 10^{5}-2.08\times 10^{6}$
$\rm Fe_{\rm 3}\rm O_{\rm 4}$ $6.1\times 10^{5}$ 5.17 0.72 $6.55\times 10^{3}-1.31\times 10^{4}$ $1.74\times 10^{4}-3.93\times 10^{4}$ $8.73\times 10^{4}-2.18\times 10^{5}$
$\gamma-\rm Fe_{\rm 2}\rm O_{\rm 3}$ $1.9\times 10^{5}$ 4.9 0.69 $1.91\times 10^{4}-3.82\times 10^{4}$ $5.09\times 10^{4}-1.14\times 10^{5}$ $2.54\times 10^{5}-6.36\times 10^{5}$
@@
Figure 14: Effect of $p$ the magnetic moment per iron atom on the relation of $p-N_{\rm H_{2}}/N_{\rm H_{2},max}$, considering model RATA--INELASTIC of SPM grains with $N_{\rm cl}=10^{3},\mu Q=3\times 10^{9}\,{\rm{erg}}\,{\rm{cm}}^{-3},a_{\rm max}=50\mu m$. Results are not filtered with CASA. By reducing $p$ from 5.5 (our adopted value) to $p=1$ (for $\rm Fe_{\rm 3}\rm O_{\rm 4}$, Yang 2021), the polarization degree reduces $\sim 2$ times, giving $p\sim 3\%$ in the inner envelope and $p<0.3\%$ in the disk scale. SPM grains must have higher $N_{\rm cl}$ than our estimation in Section 7.2 to reproduce ALMA dust polarization of $>5\%$ in the inner envelope and $\sim 1\%$ inside the disk scale.
Considering the age of our YSO of $\tau_{\rm YSOs}=38.52$ kyr (Section 2) [^13] Note that the typical timescale for Class 0 YSOs is $\sim 10^{4}$ yrs and for Class I YSOs if $\sim 10^{5}$ yrs. Given longer considered timescales, bigger iron clusters can have more time to reorient their magnetic moment to the applied magnetic field direction, which slightly increases the boundary for defining SPM., the maximum cluster volume that Brownian-like motions of clusters can overcome the anisotropy barrier energy to produce the net grain magnetic alignment is given by:
@@
One can see that for both metallic iron, $\rm Fe_{\rm 3}\rm O_{\rm 4}$, and $\gamma-\rm Fe_{\rm 2}\rm O_{\rm 3}$, adopting $N_{\rm cl}=10^{4}$ in the rest of the paper still be under the allowance range for SPM grains regardless of their position inside YSOs (except the case of $\rm Fe_{\rm 3}\rm O_{\rm 4}$ inside the envelope with $T_{\rm d}<15$ K-the case of $L_{\rm star}=5L_{\odot}$). Theoretically, we can increase $N_{\rm cl}$ to nearly $10^{5}$ to enhance the grain magnetic alignment of SPM grains inside the disk. However, given the unclear form of iron inclusions and mechanisms producing the big iron cluster inside large grains, it is very risky to adopt $N_{\rm cl}\sim 10^{5}$ to study the grain alignment dynamic inside protostellar/protoplanetary disk environments. Another related issue of $N_{\rm cr}$ is that if iron clusters are in form of $\rm Fe_{\rm 3}\rm O_{\rm 4}$ or $\gamma-\rm Fe_{\rm 2}\rm O_{\rm 3}$, dust grains which only hold big iron clusters with size $N_{\rm cl}>5\times 10^{3}$ may become ferromagnetic material inside the envelope due to the low dust temperature there (see column 5 of Table LABEL:tab:Ncr). However, ferromagnetic grains still respond with magnetic fields stronger than PM grains. Since micron-sized PM grains still can have the magnetic alignment by RATs there (Figure 2), ferro grains can also couple with $\,\boldsymbol{B}\,$ and produce polarized dust emission in the envelope scale. Indeed, as discussed in Section 7.3.3, all dust grains in the envelope do not only hold a single iron inclusion size as our assumption, and the grain can contain both big iron clusters (with $N_{\rm cl}$ may exceed $N_{\rm cr}$) and smaller ones (Yang 2021). Even big iron clusters cannot have a well-defined magnetic moment direction; the large amount of remaining small iron inclusions embedded inside dust grains do, which allows grains to still behave as SPM. As shown in Figure 8, SPM grains with $N_{\rm cl}\sim 5-10^{3}$ can already reproduce $p\sim 5-30\%$ of dust polarization inside the envelope. Consequently, even grains hold big iron clusters with $N_{\rm cl}>5\times 10^{3}$ as grains inside the disk, they can still aligned with magnetic fields, and their dust emission still helps to trace magnetic fields inside the envelope (Giang et al. 2023).
Besides the uncertainty of the upper limit of iron cluster sizes, the magnetic moment of iron atom $p$ also affects the alignment efficiency of dust grains and the observed polarization fraction. In our study, we adopt $p=5.5$ for silicate of $\rm MgFeSiO_{\rm 4}$ grains (Draine 1996, Hoang & Lazarian 2016) for calculations of SPM grains. The exact values of $p$ for iron inclusion inside protostellar grains are still unclear, that Draine (1996) suggested it to be $p=2.2-3$ (for bulk iron), Coey (2010) and Yang (2021) suggested $\sim 1.25-1.5$ (for $\rm Fe_{\rm 3}\rm O_{\rm 4}$), and $\sim 1.18-1.33$ (for $\gamma-\rm Fe_{\rm 2}\rm O_{\rm 3}$). We provide Figure 14 the $p-N_{\rm H}/N_{\rm H,max}$ relation got from different values of $p=1,2,3$, and 5.5 (our adopted value), considering model RATA--INELASTIC of SPM grains with $N_{\rm cl}=10^{3}$. The results are not filtered with CASA. One can see that by reducing $p$ by 5.5 times, the polarization degree is reduced $\sim 2$ times, similar as when we reduce $N_{\rm cl}$ by 10 times from $N_{\rm cl}=10^{3}$ to $10^{2}$ (Figure 8). The uncertainty of $p$ does not affect the dust polarization properties in the envelope, but if iron inclusions inside protostellar grains have lower $p\sim 1$, the required range of $N_{\rm cl}$ to reproduce ALMA dust polarization must be larger, i.e., the mean $N_{\rm cl}$ starting from $\geq 5\times 10^{2}$ instead of $\geq 10^{2}$ (Section 7.2) for grains in the inner envelope. Lower $p$ also further challenges micron-sized and VLGs to have magnetic alignment and produce $\sim 1\%$ of polarized dust emission inside the disk scale. As the form and the magnetic moment of iron inclusions locked inside protostellar grains remain unconstrained, these issues should be considered in further studies aiming to digest the alignment dynamic of dust grains within 100 au scale around the protostar.
@@ C.1 Whether higher $\phi_{\rm sp}$ helps to enhance the grain alignment inside the disk?
Appendix D Effect of stellar luminosity on the grain alignment state
@@
We provide the distribution map of grain alignment size in case $a_{\rm max}=10\,{\mu\rm{m}}$ in Figure 18. Similar to Figure 18, almost aligned dust grains inside protostars with low $L_{\rm center}=5L_{\odot}$ still can achieve efficient IA and magnetic alignment as grains inside objects with high $L_{\rm center}=100L_{\odot}$. But given smaller $a_{\rm max}=10\,{\mu\rm{m}}$, the narrower of the alignment size range with decreasing $L_{\rm center}$ (from $[\sim 0.8,10\,{\mu\rm{m}}]$ for $L_{\rm center}=100L_{\odot}$ to $[\sim 3-10\,{\mu\rm{m}}]$ for $L_{\rm center}=5L_{\odot}$) becomes clearer, inducing lower net observed polarization fraction shown in Figure 11, right panel).